# Detect, Investigate, Judge and Determine: A Novel LLM-based Framework for Few-shot Fake News Detection

### v1

@@ 4.4 Determination Module
More specifically, if the two predictions are identical (i.e., $P_{i}=P_{o}$), we can directly derive the final prediction with high confidence. Nevertheless, if two results diverge, indicating a conflict between the Inside Judge and Outside Judge, we further propose a determination selector to make a choice based on both sets of predictions and explanations:
$P_{f}\ =\ \mathcal{F}(Ins_{d}\uplus\ x_{test}\ \uplus$ $\ P_{i}\ \uplus R_{i}$ (6)
$\uplus\ P_{o}\ \uplus\ R_{o}\ ),$
where $Ins_{d}$ denotes the determination instruction (see Appendix A.4 for more details). $P_{f}$ is the final inference result of the DAFND model.
@@ 5.1 Experiment Setup
Datasets and Evaluation Metrics.
@@
⑤ Zephyr 60.00 63.50 66.50 48.72 53.50 54.42
⑥ ChatGLM-3 68.50 68.50 72.50 58.28 58.82 64.05
⑦ LLama-3 69.50 70.50 69.00 63.91 65.09 64.00
⑧ GPT-3.5 71.00 69.50 73.00 60.27 60.65 64.47
⑨ ARG 74.00 78.50 82.50 67.16 68.61 80.61
DAFND (ours) 86.50 87.50 88.50 81.40 82.80 84.40
@@
⑤ Zephyr 67.21 65.85 67.23 27.05 27.43 27.67
⑥ ChatGLM-3 62.49 62.75 63.43 31.59 34.83 34.15
⑦ LLama-3 65.96 65.85 66.17 30.89 35.07 31.74
⑧ GPT-3.5 68.50 69.44 67.44 32.90 36.73 36.58
⑨ ARG 61.41 77.42 76.50 42.32 51.46 46.57
DAFND (ours) 82.80 82.80 83.30 53.50 55.00 56.50
Table 2: Experimental results of our proposed method on the PolitiFact and Gossipcop datasets. Bold font represents the optimal result. For baseline methods, we follow their publicly released codes to obtain the results.
Implementation Details.
In DAFND architecture, we utilize the zephyr-7b-beta (Tunstall et al., 2023) model on Huggingface as the LLM to conduct experiments. When running Zephyr, we adhere to the default parameter values provided by the official, where the sampling temperature is $0.70$, top_k is $50$, and top_p is $0.95$. The max_new_token is set to $256$, and do_sample is set as $True$.
In the Detection Module (Section 4.1), the number of keywords to extract is set to $N=5$. In the Inside Investigation part (Section 4.2), we employ the DeBERTa-base model (He et al., 2021) from Transformers (Wolf et al., 2020) as the representation model. The number of retrieved positive/negative nearest neighbors is set as $k=2$.
All experiments are conducted on a Linux server with two 3.00GHz Intel Xeon Gold 5317 CPUs and two Tesla A100 GPUs.
Benchmark Methods.
In order to verify the effectiveness of our DAFND model, we compare DAFND with the state-of-the-art few-shot fake news detection methods. According to the model architecture, they can be grouped into three categories, including traditional fake news detection methods (① $\sim$ ④), LLM-based methods (⑤ $\sim$ ⑧), and hybrid methods (⑨).
  • PROPANEWS (Huang et al., 2023) proposes a novel framework for generating more valuable training examples, which is more beneficial to human-written situations [^5] Specifically, We combine the few-shot training set and the augmented data to train a RoBERTA-Large-based classifier, adhering to the setting in the original paper..
@@
  • MDFEND (Nan et al., 2021) incorporates the domain information through a domain gate mechanism to aggregate multiple representations extracted by a mixture of experts.
  • PSM (Ni et al., 2020) proposes to utilize Propensity Score Matching (PSM) to select decounfounded features, thereby boosting the detection of fake news.
@@
  • ChatGLM-3 (Zeng et al., 2022; Du et al., 2022) is a series of pre-trained dialogue models, and we select the ChatGLM3-6B version for the experimental comparison.
  • LLama-3 (AI@Meta, 2024) refers to the LLM proposed by Meta. We adopt its 8B version (Meta-Llama-3-8B-Instruct) sourced from Huggingface for experiments.
  • GPT-3.5 (Ouyang et al., 2022) is an advanced LLM developed by OpenAI. We leverage the API (version: gpt-3.5-turbo-0613) to conduct in-context learning.
  • ARG (Hu et al., 2024) designs an adaptive rationale guidance network for fake news detection, which integrates insights from both LLMs and traditional detection methods.
Figure 3: Ablation experiments on the PolitiFact and Gossipcop dataset.
It is worth noting that, for these LLM-based baselines (⑤ $\sim$ ⑧), we adhere to the instruction prompt proposed by (Hu et al., 2024) to conduct in-context learning. Besides, due to the limitations of maximum tokens, we randomly select 6 samples as the demonstrations, which is more than the demonstration samples utilized in our DAFND model [^6] As introduced in Section 5.1, DAFND utilizes $k=2$ positive and negative samples as demonstrations, with a total number of 4., facilitating a fair comparison [^7] For LLM-based baselines, if the LLM fails to make an inference or we are unable to categorize the output as Real or Fake categories, we directly treat it as the wrong prediction..
@@ 5.2 Experimental Result
The main results, presented in Table 2, indicate that our proposed DAFND model surpasses all baselines across various metrics, encompassing traditional, LLM-based and hybrid methods. This underscores the effectiveness of our design and the advantages of enhancing the LLM through both inside and outside perspectives. Furthermore, several notable phenomena emerge from these results:
@@
Firstly, for most baselines and our DAFND model, the performance on the PolitiFact dataset exceeds that on the Gossipcop dataset, suggesting that Gossipcop presents greater difficulty. Specifically, PolitiFact consists of political news while Gossipcop pertains to the entertainment domain. This disparity is reasonable as political news typically exhibits more organized format and content, which facilitates the fake news detection process. Secondly, with the increase of training instances ($K$), most traditional fake news detection methods (e.g., ④ PSM) and hybrid methods (⑨ ARG) show improved performance. This is logical as more data enables better training of a supervised model, mitigating the lack of prior knowledge. However, an exception is observed in ① PROPANEWS, whose performance appears relatively unaffected by $K$. As introduced in Section 5.1, different from other methods, PROPANEWS designs a data constructing strategy to supplement original training set. This significantly offsets the impact of training data quantity. Moreover, for LLM-based methods (⑤ $\sim$ ⑧), as outlined in Section 5.1, due to the limitation of maximum tokens, we all randomly select 6 samples as demonstrations. Hence, increasing the number of training instances does not substantially benefit the in-context learning of LLMs. Thirdly, although our DAFND model also faces the constraint of maximum token limitations, the $k$NN retrieval mechanism in the Inside Investigation part (Section 4.2) enhances the utilization of increased training data, thereby achieving a certain degree of improvements with higher $K$ values. Fourthly, the hybrid method (⑨ ARG), benefiting from the joint modeling of traditional models and LLMs, obtains competitive performance. And compared to that, DAFND still maintains a significant advantage, particularly in scenarios with scarcer data. These observations further demonstrate the effectiveness of our designs from multiple perspectives.
Figure 4: The case study of the DAFND model. Specifically, (a) is from the PolitiFact dataset ($K$=100), while (b) and (c) are from the Gossipcop dataset ($K$=100).
5.3 Ablation Study
@@
In this subsection, we conduct ablation experiments to assess the effectiveness of different components within our model. Specifically, we directly remove the Determination Module, consequently leading to two ablated variants: Inside Judge and Outside Judge. The results are depicted in Figure 3.
From this figure, across all configurations on PolitiFact and Gossipcop datasets, noticeable decreases are observed between the full DAFND model and its two ablated variants. This thoroughly substantiates the essentiality and non-redundancy of our DAFND designs. Notably, when the training instances are relatively limited ($K=8$), Inside Judge exhibits less effectiveness compared to Outside Judge. With the increase of training instances ($K=32,100$), Inside Judge progressively achieves more competitive performance than Outside Judge. This can be attributed to the increasing benefit of additional training data for the Inside Investigation and Inside Judge components. In contrast, Outside Investigation and Outside Judge rely on the retrieved information online, which remains unaffected by the quantity of training data.
Furthermore, considering that the backbone of DAFND is ⑤ Zephyr (Tunstall et al., 2023), we can jointly compare the results in Figure 3 and the “⑤ Zephyr” line in Table 2. This comparison highlights that both Inside Judge and Outside Judge significantly outperform Zephyr, thereby demonstrating the efficacy of our motivation to design DAFND from both inside and outside perspectives.
@@ 5.4 Case Study
To further illustrate the effectiveness of different modules in our model, we conduct a case study on both PolitiFact and Gossipcop datasets. Specifically, Figure 4 presents the input information (i.e., target news), the ground truth label, DAFND results (including the intermediate results of each module and the final results).
As depicted Figure 4 (a), the Detection Module accurately identifies the key information within the target news. And based on the valid results from Inside Investigation and Outside Investigation, both the Inside Judge and Outside Judge correctly infer that: “[This is fake news].”. These are fed into the Determination Module, leading to the final prediction, which is consistent with the ground truth label (i.e., [Fake]). Moreover, In Figure  4 (b), Inside Judge obtains the right inference (i.e., [Fake]), while Outside Judge makes the wrong inference (i.e., [Real]). Conversely, in Figure 4 (c), Inside Judge incorrectly infers “[Real]”, while Outside Judge accurately predicts “[Fake]”. In both (b) and (c), with the design of Determination Module, DAFND finally makes the correct decision. These cases intuitively demonstrate the significant role of each module in DAFND, affirming its efficacy.
More experimental analyses, such as Bad Case Analysis, can be found in Appendix B.
@@ 6 Conclusions
In this paper, we explored a motivated direction for few-shot fake news detection. We began by analyzing the limitations of current LLM-based detection methods, identifying two primary challenges: (1) Understanding Ambiguity and (2) Information Scarcity. To address these issues, we developed a Dual-perspective Augmented Fake News Detection (DAFND) model. In DAFND, a Detection Module was designed to identify keywords from the given news. Then, we proposed an Investigation Module, a Judge Module to retrieve valuable information and further generate respective predictions. More importantly, a Determination Module integrated the predictions from both inside and outside perspectives to produce the final output. Finally, extensive experiments on two publicly available datasets demonstrated the effectiveness of our proposed method. We hope our work will lead to more future studies.
Limitations
@@
In our proposed DAFND method, we need to integrate the LLM (i.e., zephyr-7b-beta introduced in Section 5.1). Due to the large scale of LLMs, it tends to consume more computing resources and time compared to traditional baselines, such as PSM (Ni et al., 2020) and FakeFlow (Ghanem et al., 2021). Essentially speaking, LLMs contain a vast amount of knowledge, much of which may be unnecessary for fake news detection. Distilling useful knowledge so as to accelerate the inference remains a valuable and intriguing research direction.
@@
Another limitation is that our current approach only employs LLMs for inference. Although we design precise prompts to implement in-context learning, it still cannot fully exploit the capabilities of LLMs due to the inherent gap between the natural language and the knowledge encoded in the model parameters. In future work, we would like to explore the low resource scenario fine-tuning techniques (e.g., lora (Hu et al., 2021a)) to adapt LLMs for the few-shot fake news detection task.
Appendix A Prompts

### v2

@@ 4.4 Determination Module
4.4.1 Inside Determination
After obtaining the effective demonstrations $\mathcal{II^{\prime}}$ from Inside Judge, we design prompts to provide the essential information to the LLM, thereby generating the inside prediction. Specifically, we first describe the target of the fake news detection task. Then, the selected inside investigation information =$\mathcal{II^{\prime}}=\{U_{positive}^{\prime},U_{negative}^{\prime}\}$ of current candidate news are followed, which augment the LLM’s understanding of this task. Finally, we prompt the LLM to predict the result of current news and give its corresponding supportive explanation:
$P_{i},R_{i}=\mathcal{F}\ (P_{ID}\ \uplus\ \mathcal{II^{\prime}}\ \uplus\ S),$ (8)
where $P_{ID}$ is the prompt instruction, $P_{i}$ refers to the prediction result, while $R_{i}$ denotes the corresponding explanation. You can move to Appendix C.2 for more details about this prompt.
4.4.2 Outside Determination
Meanwhile, with the selected outside investigation information $\mathcal{OI^{\prime}}$, we can derive the outside prediction, which is crucial for real-time news detection.
Similar to the design of Inside Judge, we describe the objective of fake news detection through an outside instruction, followed by the candidate news to be detected and the retrieved outside investigation documents $\mathcal{OI^{\prime}}$. After that, we can derive the outside prediction $P_{o}$ and explanation $R_{o}$:
$P_{o},R_{o}=\mathcal{F}\ (P_{OD}\ \uplus\ S\ \uplus\ \mathcal{OI^{\prime}}),$ (9)
4.4.3 Integrated Determination
@@
More specifically, if the two predictions are identical (i.e., $P_{i}=P_{o}$), we can directly derive the final prediction with high confidence. Nevertheless, if two results diverge, indicating a conflict between the Inside Determination and Outside Determination, we further propose a integrated selector to make a choice based on both sets of predictions and explanations:
$P_{f}=\mathcal{F}\ (P_{FD}\uplus S\uplus P_{i}\uplus R_{i}\uplus P_{o}\uplus R_{o}),$ (10)
where $P_{f}$ is the final inference result.
@@ 5.1 Experiment Setup
5.1.1 Datasets and Evaluation Metrics.
@@
⑤ KPL 58.33 73.44 82.29 60.40 73.58 81.11
⑥ Auto-CoT 49.50 58.00 64.00 53.88 58.00 55.00
⑦ Zephyr 60.00 63.50 66.50 48.72 53.50 54.42
⑧ ChatGLM-3 68.50 68.50 72.50 58.28 58.82 64.05
⑨ LLama-3 69.50 70.50 69.00 63.91 65.09 64.00
10 GPT-3.5 71.00 69.50 73.00 60.27 60.65 64.47
11 ARG 74.00 78.50 82.50 67.16 68.61 80.61
DKFND (ours) 87.00 88.00 89.00 82.43 83.78 85.33
@@
⑤ KPL 42.71 51.58 60.54 42.08 47.82 52.53
⑥ Auto-CoT 52.44 45.54 48.73 28.46 34.72 33.46
⑦ Zephyr 67.21 65.85 67.23 27.05 27.43 27.67
⑧ ChatGLM-3 62.49 62.75 63.43 31.59 34.83 34.15
⑨ LLama-3 65.96 65.85 66.17 30.89 35.07 31.74
10 GPT-3.5 68.50 69.44 67.44 32.90 36.73 36.58
11 ARG 61.41 77.42 76.50 42.32 51.46 46.57
DKFND (ours) 82.37 82.18 82.56 55.22 55.17 56.78
Table 2: Experimental results of our proposed method on the PolitiFact and Gossipcop datasets.
5.1.2 Implementation Details.
In DKFND architecture, we utilize the zephyr-7b-beta (Tunstall et al., 2023) model on Huggingface as the LLM. When running Zephyr, we adhere to the default parameter values provided by the official, where the sampling temperature is $0.70$, top_k is $50$, and top_p is $0.95$. The max_new_token is set to $256$, and do_sample is set as $True$.
In the Detection Module (Section 4.1), the number of keywords to extract is set to $N=5$.
In the Inside Investigation part (Section 4.2.1), we employ the DeBERTa-base model (He et al., 2021) from Transformers (Wolf et al., 2020) as the representation model. The number of retrieved positive/negative nearest neighbors is set as $k=5$.
In the Outside Investigation part (Section 4.2.2), we set the number of retrieved documents as $L=8$.
In the Inside Judge part (Section 4.3.1), the number of positive/negative samples is set as $m=2$.
In the Outside Judge part (Section 4.3.2), we set the number of selected documents as $n=2$.
All experiments are conducted on a Linux server with two Tesla A100 GPUs.
5.1.3 Benchmark Methods.
We compare DKFND with the state-of-the-art few-shot fake news detection methods. According to the model architecture, they can be grouped into three categories, including traditional fake news detection methods (① $\sim$ ⑤), LLM-based methods (⑥ $\sim$ 10), and hybrid methods ( 11).
  • PROPANEWS (Huang et al., 2023) proposes a framework for generating valuable training examples, beneficial to human-written situations.
@@
  • MDFEND (Nan et al., 2021) incorporates the domain information through a gate mechanism to aggregate multiple representations.
  • PSM (Ni et al., 2020) utilizes Propensity Score Matching to select decounfounded features, boosting detection performance.
  • KPL (Jiang et al., 2022) incorporates external knowledge into the prompt representation process, achieving knowledgeable detection.
  • AutoCoT (Zhang et al., 2022) proposes an automatic chain-of-thought prompting method to construct demonstrations and reasoning chains.
@@
  • ChatGLM-3 (Zeng et al., 2022; Du et al., 2022) is a series of pre-trained dialogue models, and we select the ChatGLM3-6B version.
  • LLama-3 (AI@Meta, 2024) refers to the LLM proposed by Meta. We adopt 8B version (Meta-Llama-3-8B-Instruct) for experiments.
  • 10 GPT-3.5 (Ouyang et al., 2022) is an advanced LLM developed by OpenAI. We leverage the API (version: gpt-3.5-turbo-0613) for ICL.
  • 11 ARG (Hu et al., 2024) designs an adaptive rationale guidance network, integrating insights from both LLMs and traditional methods.
It is worth noting that, for these LLM-based baselines (⑦ $\sim$ 10), we adhere to the instruction prompt proposed by (Hu et al., 2024) to conduct in-context learning. Besides, due to the limitations of maximum tokens, we randomly select 6 samples as the demonstrations, which is more than the demonstration samples utilized in our DKFND model [^3] As introduced in Section 5.1.2, DKFND utilizes $n=2$ positive and negative samples as demonstrations, with a total number of 4., facilitating a fair comparison.
@@ 5.2 Experimental Result
The main results, presented in Table 2, indicate that our proposed DKFND model surpasses all baselines across various metrics, encompassing traditional, LLM-based and hybrid methods. This underscores the effectiveness of our design and the advantages of enhancing the LLM through both inside and outside perspectives. Furthermore, several notable phenomena emerge from these results:
@@
Firstly, for most baselines and our DKFND model, the performance on the PolitiFact dataset exceeds that on the Gossipcop dataset, suggesting that Gossipcop presents greater difficulty. Specifically, PolitiFact consists of political news while Gossipcop pertains to the entertainment domain. This disparity is reasonable as political news typically exhibits more organized format and content, which facilitates the fake news detection process. Secondly, with the increase of training instances ($K$), most traditional fake news detection methods (e.g., ④ PSM) and hybrid methods ( 11 ARG) show improved performance. This is logical as more data enables better training of a supervised model, mitigating the lack of prior knowledge. However, an exception is observed in ① PROPANEWS, whose performance appears relatively unaffected by $K$. As introduced in Section 5.1.2, different from other methods, PROPANEWS designs a data constructing strategy to supplement original training set. This significantly offsets the impact of training data quantity. Moreover, for most LLM-based methods (⑦ $\sim$ 10), as outlined in Section 5.1.2, due to the limitation of maximum tokens, we all randomly select 6 samples as demonstrations. Hence, increasing the number of training instances does not substantially benefit the in-context learning of LLMs. Thirdly, although our DKFND model also faces the constraint of maximum token limitations, the $k$NN retrieval mechanism and inside judge designs enhance the utilization of increased training data, thereby achieving a certain degree of improvements with higher $K$ values. Fourthly, the hybrid method ( 11 ARG), benefiting from the joint modeling of traditional models and LLMs, obtains competitive performance. And compared to that, DKFND still maintains a significant advantage, particularly in scenarios with scarcer data. These observations further demonstrate the effectiveness of our designs from multiple perspectives.
5.3 Component Effectiveness Analysis
@@
In this subsection, we conduct ablation experiments to assess the effectiveness of different components within our model. Specifically, we first simplify the Integrated Determination part (Section 4.4.1), removing the detailed explanation in promot $P_{FD}$ (Eq.(10)). On the basis of this, we design ablated variants from inside and outside perspectives.
  • DKFND-Inside: simplifying the design of Section 4.2.1 and 4.3.1 by randomly selecting samples from training set as demonstrations.
  • DKFND-Outside: simplifying the design of Section 4.3.2 by randomly selecting retrieved outside information.
  • DKFND-Both: from both inside and outside perspectives, applying the above simplified processes to DKFND.
Figure 3: Ablation experiments on PolitiFact ($K$=100).
The results on the PolitiFact ($K$=100) dataset are depicted in Figure 3. DKFND-Inside and DKFND-Outside both exhibit a significant performance drop compared to the full DKFND. Moreover, DAFND-Both, which applies simplifications from both inside and outside perspectives, demonstrates an even greater decline in detection performance. This observation highlights the effectiveness of our proposed dual-perspective design, particularly emphasizing the irreplaceable role of information selection and decision-making designs in DKFND.
Furthermore, considering that the backbone of DKFND is ⑦ Zephyr (Tunstall et al., 2023), we can jointly compare the results in Figure 3 and the “⑦ Zephyr” line in Table 2. This comparison further reveals that all three ablation variants outperform Zephyr, demonstrating that even in its simplified forms, leveraging both internal and external perspectives for fake news detection yields significant performance gains. This finding further validates the motivation behind our dual-perspective design.
@@ 5.4 Case Study
To further illustrate the effectiveness of different modules in our model, we conduct a case study on the PolitiFact dataset. Specifically, Figure 4 presents the input information (i.e., target news), the ground truth label, DKFND results (including the inside, outside and final results).
As shown in Figure 4 (a), the Detection Module accurately identifies knowledge concepts relevant to the target news from KGs. Based on this, both Inside Determination and Outside Determination predict “[This is real news]”. They are then fed into the Integrated Determination, which correctly predicts the label as [Real], consistent with the ground truth. In Figure 4 (b), the Inside Determination incorrectly predicts [Fake], while the Outside Determination correctly predicts [Real]. Conversely, in Figure Figure 4 (c), the Inside Determination produces the correct prediction [Real], whereas the Outside Determination incorrectly classifies the news as [Fake]. In both cases, thanks to the Integrated Determination, DKFND successfully reconciles conflicting predictions and ultimately arrives at the correct classification.
More experimental analyses, such as Bad Case Analysis, can be found in Appendix AC.
@@ 6 Conclusions
In this paper, we explored a motivated direction for few-shot fake news detection. We began by analyzing the limitations of current LLM-based detection methods, identifying two primary challenges: (1) Understanding Ambiguity and (2) Information Scarcity. To address these issues, we developed a Dual-perspective Knowledge-guided Fake News Detection (DKFND) model. In DKFND, a Detection Module was designed to identify knowledge concepts from given news. Then, we proposed an Investigation Module, a Judge Module to retrieve and select valuable information. Importantly, a Determination Module integrated predictions from both inside and outside perspectives to produce the final output. Finally, extensive experiments demonstrated the effectiveness of our method.
7 Limitation
@@
In our proposed DKFND method, we need to integrate the LLM (i.e., zephyr-7b-beta introduced in Section 5.1). Due to the large scale of LLMs, it tends to consume more computing resources and time compared to traditional baselines, such as PSM (Ni et al., 2020) and FakeFlow (Ghanem et al., 2021). Essentially speaking, LLMs contain a vast amount of knowledge, much of which may be unnecessary for fake news detection. Distilling useful knowledge so as to accelerate the inference remains a valuable and intriguing research direction.
@@
Another limitation is that our current approach only employs LLMs for inference. Although we design precise prompts to implement in-context learning, it still cannot fully exploit the capabilities of LLMs due to the inherent gap between the natural language and the knowledge encoded in the model parameters. This also results in the bad case we analyzed in Section A. In future work, we would like to explore the low resource scenario fine-tuning techniques (e.g., lora (Hu et al., 2021a)) to adapt LLMs for few-shot fake news detection.
Appendix A Bad Case Analysis
In this section, we illustrate the bad cases that DKFND struggles with, with a goal to analyze its shortcomings and possible improvement directions.
As illustrated in Figure 5, a particular failure occurs. From the explanation from Inside Determination and Outside Determination modules, they both classified it into Real category because they haven’t investigated clear negative information in contrast to the given news. More specifically, this bad case reveals two potential directions for improving our model:
Figure 5: The bad case of DKFND on the Gossipcop dataset ($K$=100).
(1) For one thing, as introduced in Section 1, DKFND retrieves similar demonstrations from the training set in response to the Understanding Ambiguity problem. This relies on the assumption that valuable samples can be found in the training set, which is not always the case. As a consequence, the failure in Figure 5 occurs. Therefore, how to discriminate and mitigate the impact of such circumstances is a crucial direction for improving the DKFND design. (2) For another, as introduced in Section 4, we only employ the LLM to conduct inference, which cannot fully exploit LLMs’ powerful capabilities. In some cases, although Outside Investigation retrieves valuable information, the prediction of Outside Judge still goes wrong. In fact, despite the great reasoning ability of these general LLMs, they are not competent in news-related domains and are not sufficiently familiar with the specific expression characteristics. Thence, we would like to adopt fine-tuning techniques to adapt LLMs for the news corpus in future, which we believe could bring a positive effect to our model.
Appendix B Outside Investigatation Illustration
For better illustrate the search process of the outside investigation, we give an example in Figure 6. You can also refer to our released anonymous codes for more technical details.
Figure 6: Illustration of the Google search process.
Appendix C Prompts